40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4619     Entry time: Tue May 3 18:25:38 2011
Author: kiwamu 
Type: Summary 
Category: LSC 
Subject: PRMI locking : plan 

[Rana/Jamie/Kiwamu]

 Since we've got the PRMI locked we now should be able to do more qualitative measurements.

Here is a task list that we will measure/develop in the PRMI condition.

 

 

 - Optical gain measurements

 - Characterization of control loops

 - MICH and PRC calibrations

 - Noise budget

 - Development of automatic noise budget scripts

 - Arm loss measurement 

 - Shnupp asymmetry measurement

ELOG V3.1.3-